OpenStudy (anonymous):

lim x->infinity for ln6x/e^5x

OpenStudy (amistre64):

inf/inf invokes Lhpoital

OpenStudy (amistre64):

L'Hopital

OpenStudy (anonymous):

\[\lim x \rightarrow \infty \ln6x/e ^{5x}\]

OpenStudy (anonymous):

take the derivative of the top and bottom, don't use product rule

OpenStudy (anonymous):

0

OpenStudy (anonymous):

to apply L'H rule

OpenStudy (amistre64):

ln(6x) = ln(6) + ln(x) by the way

OpenStudy (anonymous):

are you asking this \[\lim_{x \rightarrow \infty}\frac{\ln 6x}{e^{5x}}\]

OpenStudy (anonymous):

\[e^{5x}>>\ln(6x)\]

OpenStudy (amistre64):

\[\frac{1/inf=0}{5e^{5inf}=inf}\] either way you go you get to 0

OpenStudy (amistre64):

also notice that e^5x grows alot faster than ln(6x) so it over powers it

OpenStudy (amistre64):

1/1000000000 is a good visual; which is very very tiny